# The ghost-gluon vertex in the presence of the Gribov horizon: general kinematics

### v1

@@ VII Summary and final remarks
In this work we evaluated the one-loop ghost-antighost-gluon correlation function within the RGZ framework for an arbitrary kinematic configuration and investigated the role of the running coupling. We performed the calculation for the pure gauge theory in a four dimensional Euclidean spacetime. With the purpose of comparing with other approaches, we specifically focused on three distinct kinematic configurations: the symmetric and orthogonal configurations and the soft gluon limit.
@@
This calculation is important since it serves as a benchmark to assess how well the RGZ framework describes the behavior of YM theory in the low momentum regime. In this sense this paper comes to complement previous investigations on the two-point functions at tree-level order Dudal et al. (2008, 2010) and the analysis of the ghost-gluon vertex in the case of the soft-gluon limit at one-loop order Mintz et al. (2018).
@@
The only parameter of the theory we used to fit our results to the available lattice data for the ghost-antighost-gluon vertex was the gauge coupling. All the other parameters were extracted from refs. Cucchieri et al. (2012) and Oliveira and Silva (2012), where they were determined by fitting lattice results for the gluon propagator using the RGZ tree level expression. This choice reduces the number of free parameters to the minimum, making the present analysis a more strict test of the RGZ framework.
@@
As already pointed out in Mintz et al. (2018), as for the soft gluon limit the results are compatible with lattice simulations for both SU(2) Cucchieri et al. (2008); Maas (2020) and SU(3) Ilgenfritz et al. (2007) by simply considering a constant value of the gauge coupling, i.e. independent of the momentum scale. However, this is not the case of the symmetric and orthogonal configurations, where the renormalization group effects seem to be significant. Indeed, when introducing a toy model for the running of the gauge coupling, our outcomes are quantitatively compatible with available lattice data.
@@
Our results also display qualitative agreement with other continuum approaches, specifically the dynamical DSE solutions in two different truncation schemes Aguilar et al. (2019) and the RG-improved CF model Pelaez et al. (2013); Barrios et al. (2020). Additionally, we show that quantitative differences between the RGZ results and alternative continuum approaches in the UV region can be accounted for by the particularities of the model for the RG flow of the gauge coupling.
@@
Overall, these results support the RGZ theory as a valid description of the infrared YM dynamics. There are several ways, nevertheless, in which the present analysis can be improved. A fully dynamical calculation of the two-point functions of the RGZ framework, most likely in an infrared safe scheme, would allow us to determine the running parameters of the theory by fitting the two-point functions to lattice data. Another potential study concerning this correlation function could investigate its dependence on the gauge parameter within the context of linear covariant gauges, utilizing the BRST invariant version of the RGZ action Capri et al. (2016c).
@@ A.1 Tree-level propagators
In order to calculate the ghost-antighost-gluon 3-point function in the Refined Gribov-Zwanziger theory, only a subset of the propagators of the theory are needed. These are
@@
$\langle A_{\mu}^{a}(p)\varphi_{\nu}^{bc}(-p)\rangle$ $=$ $\frac{-ig\gamma^{2}f^{abc}}{p^{4}+p^{2}(m^{2}+M^{2})+m^{2}M^{2}+2Ng^{2}\gamma^{4}}P^{\perp}_{\mu\nu}(p)$ (43)
@@ A.2 Tree-level vertices
The only vertices needed for the computation of the ghost-antighost-gluon at one-loop in the RGZ theory are
@@ Appendix B On the relation between connected and 1PI correlation functions in the presence of mixed propagators
For the present calculation of the ghost-antighost-gluon vertex, we are specifically interested in the choice
@@
$\left\langle{\bar{c}^{a}(p)\,c^{b}(q)\,A_{\mu}^{c}(k)}\right\rangle=D_{\bar{c}c}(p)D_{\bar{c}c}(q)D_{AA}(k)P_{\mu\nu}(k)\left\{\frac{\delta^{3}\Gamma}{\delta c^{a}(-p)\delta\bar{c}^{b}(-q)\delta A_{\nu}^{c}(-k)}+\frac{2g\gamma^{2}f^{cde}}{k^{2}+M^{2}}\frac{\delta^{3}\Gamma}{\delta c^{a}(-p)\delta\bar{c}^{b}(-q)\delta\varphi_{\nu}^{de}(-k)}\right\},$
@@ Appendix C Analytic expression for the ghost-gluon vertex in the soft-gluon limit
The scalar function $B_{1}(0,p)$ has been calculated (with a slightly different notation) to one-loop order in Mintz et al. (2018). The result can be written as [^6] There is a typo in the corresponding result from Mintz et al. (2018), which is corrected in this expression.

### v2

@@ VII Summary and final remarks
In this work we evaluated the one-loop ghost-gluon correlation function within the RGZ framework for an arbitrary kinematic configuration and investigated the role of the running coupling. We performed the calculation for the pure gauge theory in a four dimensional Euclidean spacetime. With the purpose of comparing with other approaches, we specifically focused on three distinct kinematic configurations: the symmetric and orthogonal configurations and the soft gluon limit.
@@
This calculation is important since it serves as a benchmark to assess how well the RGZ framework describes the behavior of YM theory in the low momentum regime. In this sense this paper comes to complement previous investigations on the two-point functions at tree-level order [43, 44] and the analysis of the ghost-gluon vertex in the case of the soft-gluon limit at one-loop order [1].
@@
The only parameter of the theory we used to fit our results to the available lattice data for the ghost-gluon vertex was the gauge coupling. All the other parameters were extracted from refs. [69] and [70], where they were determined by fitting lattice results for the gluon propagator using the RGZ tree level expression. This choice reduces the number of free parameters to the minimum, making the present analysis a more strict test of the RGZ framework.
@@
As already pointed out in [1], as for the soft gluon limit the results are compatible with lattice simulations for both SU(2) [71, 72] and SU(3) [73] by simply considering a constant value of the gauge coupling, i.e. independent of the momentum scale. However, this is not the case of the symmetric and orthogonal configurations, where the renormalization group effects seem to be significant. Indeed, when introducing a toy model for the running of the gauge coupling, our outcomes are quantitatively compatible with available lattice data.
@@
Our results also display qualitative agreement with other continuum approaches, specifically the dynamical DSE solutions in two different truncation schemes [15] and the RG-improved CF model [29, 35]. Additionally, we show that quantitative differences between the RGZ results and alternative continuum approaches in the UV region can be accounted for by the particularities of the model for the RG flow of the gauge coupling.
@@
Overall, these results support the RGZ theory as a valid description of the infrared YM dynamics. There are several ways, nevertheless, in which the present analysis can be improved. A fully dynamical calculation of the two-point functions of the RGZ framework, most likely in an infrared safe scheme, would allow us to determine the running parameters of the theory by fitting the two-point functions to lattice data. Another potential study concerning this correlation function could investigate its dependence on the gauge parameter within the context of linear covariant gauges, utilizing the BRST invariant version of the RGZ action [55].
@@ A.1 Tree-level propagators
In order to calculate the ghost-gluon vertex function in the Refined Gribov-Zwanziger theory, only a subset of the propagators of the theory are needed. These are
@@
$\langle A_{\mu}^{a}(p)\varphi_{\nu}^{bc}(-p)\rangle$ $=$ $\frac{-ig\gamma^{2}f^{abc}}{p^{4}+p^{2}(m^{2}+M^{2})+m^{2}M^{2}+2Ng^{2}\gamma^{4}}P^{\perp}_{\mu\nu}(p)=-ig\gamma^{2}f^{abc}P^{\perp}_{\mu\nu}(p)\frac{D_{AA}(p)}{p^{2}+M^{2}}$ (43)
@@ A.2 Tree-level vertices
The only vertices needed for the computation of the ghost-gluon at one-loop in the RGZ theory are
@@ Appendix B On the relation between connected and 1PI correlation functions in the presence of mixed propagators
For the present calculation of the ghost-gluon vertex, we are specifically interested in the choice
@@
$\left\langle{A_{\nu}^{a}(k)\,\bar{c}^{b}(p)\,c^{c}(q)}\right\rangle=D_{\bar{c}c}(p)D_{\bar{c}c}(q)D_{AA}(k)P_{\mu\nu}^{\perp}(k)\left\{\frac{\delta^{3}\Gamma}{\delta A_{\mu}^{a}(-k)\delta\bar{c}^{b}(-p)\delta c^{c}(-q)}-\frac{2ig\gamma^{2}f^{ade}}{k^{2}+M^{2}}\frac{\delta^{3}\Gamma}{\delta c^{b}(-p)\delta\bar{c}^{c}(-q)\delta\varphi_{\mu}^{de}(-k)}\right\},$
@@ Appendix C Analytic expression for the ghost-gluon vertex in the soft-gluon limit
The scalar function $B_{1}(0,p)$ has been calculated (with a slightly different notation) to one-loop order in [1]. The result can be written as [^8] There is a typo in the corresponding result from [1], which is corrected in this expression.